OpenStudy (anonymous):

Line AB contains points A (-1, 3) and B (5, 3). Line AB has a slope that is (5 points) Select one: a. zero b. undefined c. positive d. negative

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

m=y2-y1/x2-x1 m=3-3/5+1 m=0/6 m=0

OpenStudy (anonymous):

So its A thank you

OpenStudy (nurali):

My Pleasure.